50.00 mLof 0.1M NaOH solution is titrated with 0.05M H2SO4solution. Calculate the volume of the H2SO4 solution that is required to completely neutralize the NaOH solution. Leave your answer in 4 significant fIgures.
Answer #1

Number of moles of NaOH,n= molarity x volume in L

n= 0.1 M * 50.00 mL *10^-3 L/ mL

= 5*10^-3 mol

2 NaOH+ H2SO4 ----> Na2SO4 + H2O

2 moles of NaOH requires 1 mole of H2SO4

5*10^-3 moles of NaOH requires (5*10^-3)/2= 2.5*10^-3 moles of H2SO4

So volume of H2SO4 ,V= number of moles/ molarity

V= (2.5*10^-3 mol)/0.05 M

= 0.05 L

= 50.0 mL